# Competitive Equilibrium for Chores: from Dual Eisenberg-Gale to a Fast, Greedy, LP-based Algorithm

### v1

@@ Abstract.
We study the computation of competitive equilibrium for Fisher markets with $n$ agents and $m$ divisible chores. Prior work showed that competitive equilibria correspond to the nonzero KKT points of the Nash welfare minimization program, which is a non-convex analogue of the Eisenberg-Gale convex program. We introduce an analogue of the Eisenberg-Gale dual for chores: we show that all KKT points of this dual correspond to competitive equilibria, and while it is not a dual of the non-convex primal program in a formal sense, the objectives touch at all KKT points. Similar to the primal program, the dual has problems from an optimization perspective: there are many feasible directions where the objective tends to positive infinity, and these attract iterative optimization methods. We then derive a new constraint for the dual, which restricts optimization to a hyperplane that avoids all these directions. We show that restriction to this hyperplane retains all KKT points, and surprisingly, does not introduce any new ones. This allows, for the first time ever, application of iterative optimization methods over a convex region for computing competitive equilibria for chores.
@@
We next introduce a greedy Frank-Wolfe algorithm for optimization over our program and show a state-of-the-art convergence rate to competitive equilibrium. In the case of equal incomes, we show a $\mathcal{\tilde{O}}(n/\epsilon^{2})$ rate of convergence, which improves over the two prior state-of-the-art rates of $\mathcal{\tilde{O}}(n^{3}/\epsilon^{2})$ for an exterior-point method and $\mathcal{\tilde{O}}(nm/\epsilon^{2})$ for a combinatorial method. Moreover, our method is significantly simpler: each iteration of our method only requires solving a simple linear program. We show through numerical experiments on simulated data and a paper review bidding dataset that our method is extremely practical: it easily solves every instance we tried, usually in 5-20 iterations, is simple to implement, and has no numerical issues. This is the first highly practical method for solving competitive equilibrium for Fisher markets with chores.
conference: ; ;
1. Introduction
Competitive Equilibrium (CE) is a fundamental concept in microeconomics introduced by Arrow and Debreu (Arrow and Debreu, 1954), that studies the pricing and allocation of items to agents based on the interaction of demand and supply. Competitive Equilibrium with Equal Incomes (CEEI) is one of the prominent special cases of CE that has received substantial attention due to its remarkable applications in the fair and efficient allocation of items to the agents. In a CEEI, agents are given equal buying power. Then, the goal is to determine prices for the items, and an allocation of the items to the agents, such that (i) the total price of the bundle of items allocated to all the agents is the same (say \$$\$1$), (ii) each agent gets her most preferred bundle (utility-maximizing) of total price \$$\$1$, and (iii) all items are allocated [^1] Condition (i) distinguishes CEEI from CE in other economic models like Fisher markets and exchange markets..
The items to be allocated can be either goods giving happiness/ utility, or chores giving pain/ disutility. Throughout this paper, we assume that the agents have linear utility/ disutility functions, which is also the most extensively studied case (Goldman and Procaccia, 2015) [^2] The online platform Spliddit has seen tens of thousands of users for the past few years (Goldman and Procaccia, 2015). Spliddit uses additive preferences which is the parallel of linear utility functions in the context of dividing indivisible items. and is commonly used in applications (spl, ; Conitzer et al., 2022; Allouah et al., 2023) and online fair allocation (Azar et al., 2016; Banerjee et al., 2022; Gao et al., 2021; Benadè et al., 2023).
CE with goods.
CE with goods has seen a long line of research since the 1950s. Early work (Arrow and Debreu, 1954) showed the existence of CE, and also formulated convex programs (Eisenberg and Gale, 1959) that capture all CE under specific utility functions. In particular, (Eisenberg and Gale, 1959) show that any allocation that maximizes the product of the utility functions of the agents corresponds to a CE. Despite the existence of the clean convex program, and polynomial bounded rationality of the optimal solutions, there has been a long study on a variety of algorithms to compute CE under linear utilities, like interior point methods (Jain, 2007; Ye, 2008), combinatorial methods (Devanur et al., 2008; Orlin, 2010; Duan and Mehlhorn, 2015; Garg and Végh, 2019; Chaudhury and Mehlhorn, 2018), and first-order methods and dynamics (Birnbaum et al., 2011; Zhang, 2011; Gao and Kroer, 2020). One of the reasons for designing special-purpose algorithms is attributed to the lack of practically fast convex program solvers for large instances (Vazirani, 2012)– for instance, convex program solvers are significantly slower than LP solvers, and are practical only for small instances (see benchmarks in Mittleman). Furthermore, the broad algorithmic literature has also revealed several structural and economically valuable insights on CE (Duan et al., 2016; Devanur et al., 2013; Garg and Végh, 2019).
CE with chores.
Despite the resemblance to CE with goods, CE with chores has significantly different structural and computational properties. Firstly, the set of CE can be disconnected (Bogomolnaia et al., 2017), in contrast to CE with goods where all equilibria form a convex set [^3] All equilibria are captured by the Eisenberg-Gale convex program.. Computing CE in exchange markets with chores is PPAD-hard (Chaudhury et al., 2022), while there exists strong polynomial time algorithms in the goods setting (Garg and Végh, 2019). Polynomial-time algorithms are only known for computing approximate CEEI (Boodaghians et al., 2022; Chaudhury et al., 2022), or special cases (Brânzei and Sandomirskiy, 2023; Garg and McGlaughlin, 2020) (constant number of agents or constant number of chores).
Despite the disconnectedness of equilibria, all CE with chores can be captured by an Eisenberg Gale-like program. In the seminal paper of Bogomolnaia et al. (2017), they show that every KKT point of the program (call it the EG program for chores) that minimizes the weighted product of the disutility functions of the agents over the feasible allocation space, subject to the disutility of each agent being strictly positive, corresponds to a CE. Note that this program has an open constraint of strictly positive disutility for each agent, and the global infimum is achieved at the closure of this open constraint. Namely, when the disutility of one of the agents is made zero (by not allocating any chore to her), the weighted product-of-disutilities objective becomes zero which is the minimum possible value. Boodaghians et al. (2022) show that first-order (and higher-order) optimization methods are attracted towards the limit points where the smallest disutility is zero, thereby not converging to a desired KKT point.
Product to sum-log: Open constraints to an unbounded objective. For ease of presentation, we choose a different representation of the same problem in this paper: we change the objective from minimizing the weighted product of disutilities to minimizing the weighted sum of logarithms of the disutilities. Note that the domain of the logarithm function automatically rules out zero disutility and, thereby, the degenerate allocations from the set of feasible allocations. However, the same problem persists: there is an unbounded-objective direction, i.e., by making the disutility of an agent tend to zero, we can have the objective tend to negative infinity, and standard iterative optimization methods move along these unbounded objective directions.
To circumvent this problem, sophisticated iterative methods have been developed (Boodaghians et al., 2022; Chaudhury et al., 2022) to find a desired KKT point. However, both methods, involve solving a non-linear convex program at each iteration – the projection subroutine in (Boodaghians et al., 2022), and the balanced flow subroutine in (Chaudhury et al., 2022). Given that general-purpose non-linear convex program solvers are seldom effective for large instances in practice (Charkhgard et al., 2018), this necessitates the study of simple, practically implementable approaches that do not rely on solving non-linear programs. The first attempt towards such an algorithm would be to fix the issue of unbounded-objective directions with the existing program, i.e., to come up with a program that captures all CE as its KKT points (similar to the existing EG program for chores), and that does not have exhibit unbounded-objective directions. If we can answer the foregoing question affirmatively, the next direction of investigation would be whether there are efficient first-order methods to find the desired KKT points, and how these methods compare to existing baselines (in theory and in practice). In this paper, we answer both questions affirmatively.
1.1. Our Contributions
Novel program with no unbounded-objective direction:
@@ 1.1. Our Contributions
As our first main result, we give a novel convex maximization program subject to simple linear constraints. We show that our program circumvents the issue of unbounded-objective directionof the EG program for chores while maintaining one-to-one correspondence between CE and KKT points. To the best of our knowledge, this is the first program that makes the chores problem amenable to standard gradient-descent-type algorithmic approaches.
@@
The inspiration for our novel program comes from the dual of the EG program in the context of goods. The dual EG program for goods has been used algorithmically by Birnbaum et al. (2011) to investigate the convergence of proportional response dynamics, and by Gao et al. (2021) for computing a CEEI with goods online. So naturally, we investigate whether a similar dual formulation for CE with chores exists. We show that, indeed, there is a natural “dual” formulation of the EG program for chores. [^4] We remark that since the primal program is non-convex, this is not a dual in the strict sense. However, just as the primal EG program for chores, the natural dual EG program for chores also exhibits the problem of unbounded-objective direction. The dual program also admits additional unbounded-objective directions– in particular, unbounded directions along which we can make the objective tend to positive infinity. However, surprisingly, we show that all the unbounded-objective directions can be removed by adding an extra linear constraint. This constraint not only gets rid of all unbounded-objective directions, but also preserves all KKT points and does not introduce any new KKT point. We call the new program Chores Dual Redundant. We observe that, although a non-convex optimization problem, Chores Dual Redundant has strong duality-like structural properties: its objective matches that of the primal EG program at the KKT points. The full details of the novel program, its properties, and its connections with the primal program are discussed in Section 3.
Theorem 0 ().
Given an instance of chores allocation, there exists a one to one correspondence between CE and the KKT points of Chores Dual Redundant. Furthermore, Chores Dual Redundant exhibits no unbounded-objective directions.
Greedy Frank Wolfe Algorithm:
Armed with this new program it is natural to attempt a first-order method to compute a CEEI with chores. We adopt the greedy variant of the famous Frank-Wolfe method (Frank and Wolfe, 1956) (FW). Given a convex function $f$, the FW algorithm, in every iteration $k$, finds a minimizer $s_{k}$ to the linear approximation of the function around the previous iterate point $x_{k-1}$. Thereafter, the algorithm optimally chooses a step $\alpha_{k}$ towards $s_{k}$ from $x_{k-1}$ and sets $x_{k}=x_{k-1}+\alpha_{k}(s_{k}-x_{k-1})$ for $\alpha_{k}\in[0,1]$. The advantage of this method over other first-order methods like gradient descent is that it avoids projection (which can often be expensive). Although this algorithm is mostly used in the context of convex minimization (Jaggi, 2013), we can also adopt this algorithm for our non-convex program. In particular, our (Chores Dual Redundant) program is a convex maximization program, and it turns out that we can greedily choose our current iterate point to be the maximizer $s_{k}$ of the linear approximation of the function around $x_{k-1}$. That is, we set $x_{k}=s_{k}$ where $s_{k}={\operatorname{\mathrm{arg\,max}}}_{s\in F}\langle\nabla f(x_{k-1}),s-x_{k-1}\rangle$, where $F$ is the space of feasible points. We call our algorithm the greedy Frank Wolfe method (GFW), discussed in detail in Section 3.
We next give the reader a high-level intuition on why our GFW technique works. Since our objective function $f$ is strictly convex, we have,
$f(x_{k})-f(x_{k-1})>\langle\nabla f(x_{k-1}),x_{k}-x_{k-1}\rangle$
Note that since $s_{k}$ is the maximizer for $\langle\nabla f(x_{k-1}),x_{k}-x_{k-1}\rangle$, and $x_{k-1}$ is also a feasible point, we have
$\langle\nabla f(x_{k-1}),x_{k}-x_{k-1}\rangle$ $=\langle\nabla f(x_{k-1}),s_{k}-x_{k-1}\rangle$
$\geq\langle\nabla f(x_{k-1}),x_{k-1}-x_{k-1}\rangle=0.$
Therefore, we have a sequence of iterates that improves our objective function. Furthermore, since $\max_{s\in F}\langle\nabla f(x_{k-1}),s-x_{k-1}\rangle$ is an LP [^5] Recall that $F$ is a polytope as we have only linear constraints in Chores Dual Redundant, we can assume that each $x_{k}$ is a vertex of $F$, implying that after finitely many iterations, our objective will not improve, i.e., for all $s\in F$, we have $\langle\nabla f(x_{k-1}),s-x_{k-1}\rangle=0$, which would correspond to a KKT point of Chores Dual Redundant. In Section 4, we show that this method also converges to an approximate $\varepsilon$-approximate CEEI (formally defined in Definition 2) in $\mathcal{O}(1/\varepsilon^{2})$ iterations [^6] This hides polynomial factors in the number of agents and number of chores, matching the convergence guarantees in  (Boodaghians et al., 2022; Chaudhury et al., 2022), in terms of dependence on $\epsilon$, while improving the dependence on the number of buyers and chores. This is where the main technical bulk of the paper lies.
Theorem 0 ().
The GFW method finds a $\varepsilon$-approximate CEEI in $\mathcal{O}(n\log(mD)/\varepsilon^{2})$ iterations, where $n,m$ and $D$ represent the number of agents, number of chores and the largest disutility an agent can have for one unit of any chore.
[…]
We emphasize that unlike the algorithms in (Boodaghians et al., 2022; Chaudhury et al., 2022), every iteration of GFW involves solving only an LP ($\max_{s\in F}\langle\nabla f(x_{k-1}),s-x_{k-1}\rangle$). Further, the upper bound on the number of iterations of GFW is strictly smaller than the upper bound on the number of iterations of the algorithms in (Boodaghians et al., 2022; Chaudhury et al., 2022). Therefore, we expect it to be considerably faster in practice, especially for large instances. We confirm this intuition through a comprehensive set of empirical results.
@@
We compare our GFW method to the state-of-the-art existing method: the exterior-point method (EPM) from Boodaghians et al. (2022), on a large number of instances sampled from five different classes of random disutility distributions, as well as on a semi-synthetic dataset derived from a conference review bidding dataset with noise added. Our GFW algorithm solves every instance that we tried it on, usually in 10-30 iterations, and in less than a minute for instances with several hundred buyers and chores. In contrast, the prior EPM fails on a large number of instances (for some disutility distributions it fails on every instance once the number of buyers and chores is on the order of a few hundred each), though it is usually competitive with our method in the cases where it does not fail. We conclude that our GFW method is the first highly practical method: it is extremely robust and simple to implement, and can computes an exact CE within a minute for every instance we tried, even with upwards of 300 buyers and chores.
2. Preliminaries
Chores Market.
[…]

### v2

@@ Abstract.
We study the computation of competitive equilibrium for Fisher markets with $n$ agents and $m$ divisible chores. Competitive equilibria for chores are known to correspond to the nonzero KKT points of a program that minimizes the product of agent disutilities, which is a non-convex program whose zero points foil iterative optimization methods. We introduce a dual-like analogue of this program, and show that a simple modification to our “dual” program avoids such zero points, while retaining the correspondence between KKT points and competitive equilibria. This allows, for the first time ever, application of iterative optimization methods over a convex region for computing competitive equilibria for chores. We next introduce a greedy Frank-Wolfe algorithm for optimization over our program and show a new state-of-the-art convergence rate to competitive equilibrium. Moreover, our method is significantly simpler than prior methods: each iteration of our method only requires solving a simple linear program. We show through numerical experiments that our method is extremely practical: it easily solves every instance we tried, including instances with hundreds of agents and up to 1000 chores, usually in 10-30 iterations, is simple to implement, and has no numerical issues.
1 Introduction
We study competitive equilibrium (CE) for Fisher markets with chores, where a set of divisible chores need to be allocated among a set of agents. Each agent incurs disutility from the chores assigned to her, and is compensated with payments for the chores she undertakes. We study the most classical utility model, where the disutility of each agent is linear in the chores assigned to her. Each agent is assumed to have a fixed earning requirement. CE is a central solution concept in microeconomics which specifies how markets can assign prices and allocations in a way that satisfies demand and supply. In the case of Fisher markets with chores, prices and allocations form a CE if (i) each agent gets her disutility-minimizing bundle that achieves her earning requirement, and (ii) all chores are fully allocated.
A classical motivation for the Fisher market model with goods, is the competitive equilibrium with equal incomes (CEEI), which is a mechanism for fairly and efficiently dividing a set of goods among a set of agents (Varian 1974). The CEEI idea naturally extends to the case of chores, and still provides similar fairness and efficiency properties. Examples of fair chore allocation problems include family members dividing household chores, workers dividing job shifts, faculty dividing teaching loads, peer review assignment, and so on.
In a seminal paper, Bogomolnaia et al. (2017) showed that CE exists in chores Fisher markets with 1-homogeneous convex disutility functions [^1] Bogomolnaia et al. (2017) measure utility, in which case the utility function is concave; for our purposes it is easier to work with the equivalent disutility of an agent, in which case the corresponding function is convex., and gave an optimization characterization of the set of CE. They also showed that, despite the resemblance to CE with goods, CE with chores has surprisingly different structural and computational properties. Firstly, the set of CE can be disconnected (Bogomolnaia et al. 2017), in contrast to CE with goods where all equilibria form a convex set. Secondly, the CE set for goods is the set of optimal solutions to a convex program known as the Eisenberg-Gale convex program, which is a program that maximizes the product of agent utilities. In the case of chores, the set of CE is still captured by an analogous program which attempts to minimize the product of agent disutilities. Bogomolnaia et al. (2017) show that every KKT point of this program (call it the EG program for chores) where every agent receives strictly positive distuility corresponds to a CE. Importantly, KKT points where some agent receives zero disutility do not correspond to CE, yet such points are the global minimizes of the program. This effectively means that EG for chores has an open constraint, and the global infimum is achieved at the closure of this open constraint. This is highly problematic from an optimization perspective, because these zero points may attract iterative optimization methods. Due to this algorithmic difficulty, polynomial-time algorithms are only known for computing approximate CEEI (Boodaghians et al. 2022, Chaudhury et al. 2022), and existing methods all rely on highly specialized subroutines that are not very practical.
As is the case for the EG program for goods, we will work with the logarithm of the product of agent disutilities, which converts the product of disutilities into a concave function, albeit one that we need to minimize. Note that the domain of the logarithm function automatically rules out allocations with zero disutility for some agent, and thus all KKT points of this objective correspond to CE. However, this does not solve the problem of applying iterative optimization methods: under this formulation the objective has infeasible points on the boundary, corresponding to zero disutility allocations, such that the objective converges to negative infinity as we approach these points. We refer to these points as poles, in analogy to their usage in complex analysis. In this formulation, starting from almost every feasible points, standard iterative optimization methods move towards these poles and the objective tends to negative infinity (Boodaghians et al. 2022).
To circumvent the poles issue, sophisticated and highly specialized iterative methods have been developed (Boodaghians et al. 2022, Chaudhury et al. 2022) for finding a nonzero KKT point. However, both methods require exactly solving a non-linear convex program at each iteration – the projection subroutine in (Boodaghians et al. 2022), and the balanced flow subroutine in (Chaudhury et al. 2022). Exactly solving a nonlinear convex program at every iteration can be expensive, and thus these methods are not practical for large-scale problems. A major roadblock for developing more practical algorithms is the presence of poles in the EG for chores program. This opens up a natural question:
Does there exists an optimization formulation of the chores problem that avoids the poles issue?
An affirmative answer to this question would open up the CE for chores problem to more standard iterative optimization methods, making it more tractable both theoretically and practically. As we discuss below, we show that such a formulation indeed exists, and it leads to significantly faster algorithms.
Organization.
The remainder of this section discusses our contributions and related work. Section 2 gives preliminaries for Fisher markets with chores and existing mathematical programs. Section 3 introduces two new (dual EG) programs capturing CE with chores. Section 4 provides with a new greedy, fast, LP-based algorithm to compute a CE, and proves its convergence properties. Section 5 demonstrates practical efficiency of our new algorithm and Section 6 includes conclusion and discussion.
1.1 Our Contributions
Novel program with no poles.
@@ 1.1 Our Contributions
As our first main result, we give a novel program that maximizes a convex function subject to simple linear constraints. We show that our program circumvents the issue of poles of the EG program for chores, while maintaining the one-to-one correspondence between CE and KKT points. To the best of our knowledge, this is the first program that makes the chores problem amenable to standard gradient-descent-type algorithmic approaches.
@@
The inspiration for our new program comes from the dual of the EG program in the context of goods. The dual EG program for goods has been used extensively for deriving effective algorithms, as we discuss in Section 1.2. In a formal sense, there is no duality theory for the chores setting, since we are dealing with nonconvex programs. Nonetheless, we develop a natural “dual” formulation of the EG program for chores. However, just as the primal EG program for chores, the natural dual EG program for chores also has poles. Surprisingly, we show that the poles of our new program can be removed by adding a linear constraint. This constraint not only gets rid of all poles, but also preserves all KKT points without introducing any new KKT points. We call the new program (Chores Dual Redundant). We observe that, although a non-convex optimization problem, (Chores Dual Redundant) has strong duality-like structural properties: its objective matches that of the primal EG program at the KKT points.
Greedy Frank Wolfe algorithm.
Armed with this new program, it is natural to attempt a first-order method to compute a CE with chores. In this paper, we adopt the greedy variant of the famous Frank-Wolfe (FW) method (Frank and Wolfe 1956). In Section 4, we show that this method converges to an $\epsilon$-approximate CE (formally defined in Definition (2)) in $\mathcal{\tilde{O}}(\frac{n}{\epsilon^{2}})$ iterations, matching the convergence guarantees in Boodaghians et al. (2022), Chaudhury et al. (2022), in terms of dependence on $\epsilon$, while improving the dependence on the number of buyers and chores. To be specific, the exterior point method by Boodaghians et al. (2022) requires $\mathcal{\tilde{O}}(\frac{n^{3}}{\epsilon^{2}})$ iterations, and the combinatorial algorithm by Chaudhury et al. (2022) requires $\mathcal{\tilde{O}}(\frac{nm}{\epsilon^{2}})$ iterations. We emphasize that, unlike the algorithms in Boodaghians et al. (2022), Chaudhury et al. (2022), an iteration of GFW involves solving only a single LP, as opposed to more general convex programs in prior methods. Further, the dependence on problem size in the bound on the number of iterations for our GFW method is significantly smaller than prior methods (Boodaghians et al. 2022, Chaudhury et al. 2022). Combining these facts, we expect GFW to be considerably faster in practice, especially for large instances. We confirm this intuition through a comprehensive set of empirical results.
Convergence for relative strongly convex maximization.
Consider a problem of maximizing a relative-strongly-convex function $f$ (w.r.t. some reference function $h$) over a nonempty closed convex set. We show that after $T\in\mathcal{O}\Big{(}\frac{f^{*}-f^{0}}{\alpha\epsilon}\Big{)}$ [^2] Here, $f^{*}$ and $f^{0}$ are the optimal and initial objective values, respectively. $\alpha$ denotes the modulus of the relatively strong convexity. iterations of the GFW algorithm, there exists an iterate $k\in[T]$, such that the distance between $x_{k}$ and $x_{k-1}$ is small compared to $\epsilon$, as measured by the Bregman divergence induced by the reference function. See Section 4.1 for our generalized results. We use this result, along with a particular choice of $h$, to show that GFW finds an $\epsilon$-approximate CE. We believe that this general approach could lead to interesting guarantees for other problems of maximizing a convex function (e.g. the approximate $\ell_{0}$-penalized PCA [^3] Principal component analysis. problem (Luss and Teboulle 2013), multiplicative problems (Benson 1995), etc.). In the appendix, we also present an alternative primal-dual analysis of GFW specifically for the (Chores Dual Redundant) program, with mildly stronger guarantees (see Appendix A).
@@
We compare our GFW method to the state-of-the-art existing method: the exterior-point method (EPM) from Boodaghians et al. (2022), on a large number of instances sampled from five different classes of random disutility distributions, as well as on a semi-synthetic dataset derived from a conference review bidding dataset. Our GFW algorithm solves every instance that we tried it on, usually in 10-30 iterations, and in less than a minute for instances with several hundred buyers and chores. In contrast, the prior EPM fails on a large number of instances (for some disutility distributions it fails on every instance once the number of buyers and chores is on the order of a few hundred each), though it is usually competitive with our method in the cases where it does not fail. We conclude that our GFW method is the first highly practical method: it is extremely robust and simple to implement, and can computes an exact CE within a minute for every instance we tried, even with upwards of 300 buyers and chores.
1.2 Related Work
CE with goods has seen a long line of research since the 1950s. Early work (Arrow and Debreu 1954) showed the existence of CE, and also formulated convex programs (Eisenberg and Gale 1959) that capture all CE under specific utility functions. In particular, Eisenberg and Gale (1959) show that any allocation that maximizes the product of the utility functions of the agents corresponds to a CE. Despite the existence of the convex program, and polynomial bounded rationality of the optimal solutions, there has been a long study on a variety of algorithms to compute CE under linear utilities, including interior point methods (Jain 2007, Ye 2008), combinatorial methods (Devanur et al. 2008, Orlin 2010, Duan and Mehlhorn 2015, Garg and Végh 2019, Chaudhury and Mehlhorn 2018), and first-order methods and dynamics (Birnbaum et al. 2011, Zhang 2011, Gao and Kroer 2020). One of the reasons for designing special-purpose algorithms is attributed to the lack of practically fast convex program solvers for large instances (Vazirani 2012)– for instance, convex program solvers are significantly slower than LP solvers, and are practical only for medium-size instances (see benchmarks in Mittelmann). Furthermore, the broad algorithmic literature has also revealed several structural and economically valuable insights on CE (Duan et al. 2016, Devanur et al. 2013, Garg and Végh 2019). Finally, CE has been used in a variety of applications, especially in the case of linear utility or disutility functions (Goldman and Procaccia 2015, Conitzer et al. 2022, Allouah et al. 2023). For example, the online platform Spliddit has seen tens of thousands of users for the past few years (Goldman and Procaccia 2015). Spliddit uses additive preferences which is the parallel of linear utility functions in the context of sharing indivisible items. CE with linear utilities has also seen extensive use in online fair allocation (Azar et al. 2016, Banerjee et al. 2022, Gao et al. 2021, Benadè et al. 2023).
The problem of computing a CE involving chores has attracted significant attention since the characterization result of Bogomolnaia et al. (2017). There are polynomial-time algorithms if the number of agents or the number of chores is a constant: in this special case, Brânzei and Sandomirskiy (2023) showed CE with chores can be computed in strongly polynomial time and Garg and McGlaughlin (2020) showed a polynomial-time algorithm for computing CE in settings that include both goods and chores. Chaudhury et al. (2021) developed a simplex-like algorithm based on Lemke’s scheme for additively separable piecewise linear concave utilities, and showed that it can be implemented for small instances with 10-20 agents and chores. As mentioned previously Boodaghians et al. (2022) and Chaudhury et al. (2022) developed polynomial-time algorithms for computing approximate CE with chores, where their iteratations require solving non-linear convex programs.
2 Preliminaries
2.1 Fisher Market with Chores
Analogous to Fisher markets in the goods setting, a chores market comprises of a set of $n$ agents, and a set of $m$ divisible chores. Each agent $i$ has a disutility of $d_{ij}$ for one unit of chore $j$. Given a bundle of chores $x_{i}\in\mathbb{R}^{m}_{\geq 0}$, where $x_{ij}$ denotes the amount of chore $j$, agent $i$’s total linear disutility for bundle $x_{i}$ is $\sum_{j\in[m]}d_{ij}x_{ij}$. We also write $\sum_{j\in[m]}d_{ij}x_{ij}$ as $\langle d_{i},x_{i}\rangle$, where $d_{i}=(d_{i1},d_{i2},\dots,d_{im})$. Further, each agent $i$ has an earning requirement of $B_{i}\geq 0$ units of money.